OpenStudy (anonymous):

12In an experiment, a petri dish with a colony of bacteria is exposed to cold temperatures and then warmed again. Find a quadratic model for the data in the table. Type your answer below. Show your work. Time (hours) 0 1 2 3 4 5 6 Population (1000s) 5.1 3.03 1.72 1.17 1.38 2.35 4.08

OpenStudy (anonymous):

@amberxoxo

OpenStudy (anonymous):

sh*t

OpenStudy (anonymous):

ikr

OpenStudy (anonymous):

whys everythig gotta be math

OpenStudy (anonymous):

becuase im in alebre 2 right now lol

OpenStudy (anonymous):

oh lol ok

OpenStudy (anonymous):

i take one class at a time

OpenStudy (anonymous):

lol. math is stupid

OpenStudy (anonymous):

̿ ̿̿ ̿’̿’̵͇̿̿з=(◣_◢)=ε/̵͇̿̿/’̿’̿ ̿ ̿̿ ̿̿ kill issac newton and all those math people

OpenStudy (anonymous):

yah ikr lol

OpenStudy (anonymous):

@Lakeekee_ i have to show my work on dis doe

OpenStudy (anonymous):

ight holup let mi work dis out

OpenStudy (anonymous):

ok fam

OpenStudy (anonymous):

http://staff.bbhcsd.org/bradacm/files/2013/10/Enriched-alg-2-trig-unit-2-test-review-key.pdf this link has the work and the problem and i think it will help

OpenStudy (anonymous):

wich number is it

OpenStudy (anonymous):

24

OpenStudy (anonymous):

all this work lol thnx b

OpenStudy (anonymous):

Let the equation be : at^2 + bt + c = P, where t = time (hrs), P = population (1000's). When t = 1, P = 3.03. When t = 2, P = 1.72. When t = 3, P = 1.17. Substitute these into the equation to obtain these 3 simultaneous equations : a + b + c = 3.03 4a + 2b + c = 1.72 9a + 3b = c = 1.17 Solving gives : a = 0.38, b = -2.45, c = 5.1. The equation is therefore, P = 0.38t^2 - 2.45t + 5.1 Testing with t = 0 to 6 gives the population values as provided, so it seems to be a valid model. At t = 9 hrs, P = 0.38*9^2 - 2.45*9 + 5.1 = 13.83.

OpenStudy (anonymous):

Use the model from problem 12 to estimate the population of bacteria at 9 hours. Type your answer below. Show your work.

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

next qution for dis table lol

OpenStudy (anonymous):

what question? on the link

OpenStudy (anonymous):

opps nvr mind lol

OpenStudy (anonymous):

yhu st8 lol anymo?

OpenStudy (anonymous):

yes but give me a second to write down these mess of khouas of words that look like dis to me fsabgvrakvntkwvetivtrweitvniv3qginlwrev lol

OpenStudy (anonymous):

ight and im so dead rn lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

mane you takin 55years to write i done went to mars and back

OpenStudy (anonymous):

Evaluate the expression for the given value of the variable(s). Show your work. 4(3h-6)/1+h;h=2

OpenStudy (anonymous):

i might have fdged that up lol hold on

OpenStudy (anonymous):

ok were good

OpenStudy (anonymous):

ight hol up let me solve this one

OpenStudy (anonymous):

ight fam

OpenStudy (anonymous):

amber she is a miracle worker cuz if she didnt answer these i would have failed life lol

OpenStudy (anonymous):

4 plus 6h 10plus 2 divide by this answer go x=158 over 4-6 times 2 168times 1000 plus 4-9 total= 1038694x

OpenStudy (anonymous):

dang that took forever

OpenStudy (anonymous):

lolo ok i have five nire question you ready fam lol

OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, -20), (0, -4), (4, -20). Show your work.

OpenStudy (anonymous):

wrf is yhu takin a test boi??

OpenStudy (anonymous):

naw its my question i ahve failed over all my first semsetr lol i only have 5 left

OpenStudy (anonymous):

all ight what is dey

OpenStudy (anonymous):

i have time for this question becuase i ahve to go to work soon lol will do the rest tomorrrow lol thnx doe

OpenStudy (anonymous):

yw boo jus hit me up if ya need my help

OpenStudy (anonymous):

ok can you do that last question in two minutes plz i would greatly aprreciate very much

OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, -20), (0, -4), (4, -20). Show your work

OpenStudy (anonymous):

ight holup